Please help! I have tried for a while, but I can't come up with the correct solution for these problem types:(not from a book)
Roberto invested some money at 7%, and then invested $3000 more than twice this amount at 12%. His total annual income from the two investments was $4700. How much was invested at 12%?

Answer by checkley71(8403)

.07x+.12(x+3000)=4700
.07x+.12x+360=4700
.19x=4700-360
.19x=4340
x=4340/.19
x=$22,842 invested @7%
22,842+3000=$25,842 invested @ 12%
proof
.07*22,842+25842*.12=4700
1598.95+3101.05=4700
4700=4700